# A Unified Sequence Parallelism Approach for Long Context Generative AI

### v3

@@ Abstract
Sequence parallelism (SP), which divides the sequence dimension of input tensors across multiple computational devices, is becoming key to unlocking the long-context capabilities of generative AI models. This paper investigates the state-of-the-art SP approaches, i.e. DeepSpeed-Ulysses and Ring-Attention, and proposes a unified SP approach, which is more robust to transformer model architectures and network hardware topology. This paper compares the communication and memory cost of SP and existing parallelism, including data/tensor/zero/pipeline parallelism, and discusses the best practices for designing hybrid 4D parallelism involving SP. We achieved 86% MFU on two 8xA800 nodes using SP for sequence length 208K for the LLAMA3-8B model. Our code is publicly available on https://github.com/feifeibear/long-context-attention.
@@ 2 Sequence Parallelism Approaches
Notation Description
$L$ Sequence Length
$d$ Hidden Dimension
$hc$ Head Count
$hs$ Head Size
$bs$ Batch Size
$N$ Device Number
@@
The readiness of the standalone SP technology is marked by the publication of two milestone papers in November 2023. The DeepSpeed-Ulysses [1], named as SP-Ulysses, and Ring-Attention [2], named as SP-Ring, solving the longstanding memory and communication issues inherent to SP from two distinct perspectives. For both methodologies, each computational device is allocated a distinct segment of the $Q$ (query), $K$ (key), $V$ (value), and $O$ (output) tensors, which are segregated along the dimensions of the sequence. There is no redundancy in their storage across devices, which is a primary distinction from the early SP design [6].
@@ 3 Unified Ulysses-Ring Sequence Parallelism
The inputs to the Unified_SP_Attention include the segment of the $Q,K,andV$ tensors after being partitioned along the sequence dimension and the output is a tensor $O$ shard. The size of a tensor segment is ($bs$, $L/N$, $hs$, $hd$). Note that when using MAQ, the shape of the $hc$ for the $K$ and $V$ tensor segments differs from that of the Q tensor. During forward propagation, $scatter\_idx$ is set to 1 and $scatter\_idx$ is set to 2. The AllToAll4D merges the $L$ dimension and partitions the $hc$ dimension of the $Q$, $K$, and $V$ tensors, transforming them into ($hc/N$, $bs$, $L$, $d$). They also partition the $O$ tensor along the $L$ dimension and merge the $hc$ dimension. During the backward propagation, $scatter\_idx$ is set to 2, and $scatter\_idx$ is set to 1.
Algorithm 1 Unified Single-Process Attention Implementation
1:function Unified_SP_Attention($ulysses\_pg$, $ring\_pg$, $Q$, $K$, $V$, $scatter\_idx$, $gather\_idx$)
2:     $Q\leftarrow\textbf{AllToAll4D}{Q,scatter\_idx,gather\_idx,group=ulysses\_pg}$
3:     $K\leftarrow\textbf{AllToAll4D}{K,scatter\_idx,gather\_idx,group=ulysses\_pg}$
4:     $V\leftarrow\textbf{AllToAll4D}{V,scatter\_idx,gather\_idx,group=ulysses\_pg}$
@@
6:     $O\leftarrow\textbf{AllToAll4D}{ulysses\_pg,O,gather\_idx,scatter\_idx,group=ulysses\_pg}$
@@
Figure 3: Loading Balancing Partition for SP-Ring.
The solution to the load-unbalancing issue is to reorder the input sequence tokens along the sequence dimension, as depicted on the right side of Figure 3. In the figure, the input sequence consists of 16 tokens. Under even partition, GPU0 processes tokens 0-3, while GPU3 handles tokens 12-15. After the reorder operation for load balance partitioning, GPU0 now processes tokens 0, 1, 14, 15, and GPU3 processes tokens 5,6,11,12. The workloads handled by each GPU are perfectly balanced, which is a superior solution to SP-Ring load balance [12].
The aforementioned load balancing method is quite straightforward and similar implementation has already been applied in Megatron-LM. We present its application for Unified SP method. The Algorithm 2 extracts and reorders the input sequence from the global input sequence in hybrid Ulysses and Ring parallelism. For the positional encoding only involves element-wise operations, i.e. RoPE [13], the same reordering operation also needs to be applied to the positional encoding parameters of the model. Since the operation is applied to the model’s input token sequence, which is an integer vector of length $bs*L$, the additional overhead for load balancing is negligible.
@@
7:     $local\_seq\leftarrow reorder\_seq.chunk(ud,dim=dim)[u\_rank]$
@@
Firstly, Unified SP can remove the head number limitation of SP-Ulysses. For example, to run llama3-8B  [^3] https://github.com/meta-llama/llama3 with $hc$=8 in a 16-degree SP, one can set the Ulysses degree to 8 and the ring degree to 2.
Secondly, Unified SP allows lower bandwidth and topology requirements for network infrastructure by offering a more robust communication pattern. By setting the Ulysses degree to a value between 1 and N, the Attention communication pattern will be a mix of P2P and All2All as shown in Figure 4. Such a communication pattern is particularly well-suited for heterogeneous communication networks, allowing All2All operations to operate in high-bandwidth interconnections while asynchronous P2P communications operate in lower-bandwidth sections. This is applicable to scenarios such as a node in which GPUs connected via PCIe Switch or a cluster of GPU nodes between which the network is Ethernet connected.
@@
Figure 4: The Unified-SP is more robust to network hardware requirements.
@@ 4 SP in 4D Parallelism
In terms of communication performance, SP is inferior to DP. Both SP and DP require the allreduce operation on gradients during backward propagation. The difference in their communication performance lies in the attention module, where SP introduces additional communications overhead for activations. When the Ulysses degree is set to be greater than 1, the communication overhead of SP will be larger than that of DP due to the all2all operations. When using the Ring method, although the additional P2P communication for attention is overlapped, it introduces extra performance issues. The ideal performance is only to reach the performance of DP without communication for attention.
In terms of memory performance, both SP and DP are equivalent, as they all can reduce the activation footprint to 1/N.
@@
ZeRO [14] is a distributed parameter management method that reduces the storage space requirements of each computing device by sharding the Optimizer States (ZeRO-1), Gradients (ZeRO-2), and Parameters (ZeRO-3) across multiple devices. The memory cost for Optimizer States, Gradients, and Parameters is reduced to 1/$N$ of the original.
ZeRO can also operate within an SP process group because partitioning along the batch dimension ($bs$) or the sequence dimension ($L$) is equivalent to ZeRO’s approach. For ZeRO, SP and DP process groups can be viewed as a combined process group of size $N_{s}p\times N_{d}p$.
@@
PP partitions transformer blocks across layers, so it is complementary with the SP, which partitions tensors inside a transformer block. Therefore, SP and PP are fully compatible.
Since SP can form a unified parallel group with ZeRO for collective communication, we believe TP should still be placed at the lowest dimension of the 4D parallel group partitioning.
@@ 5.1 Performance of Unified SP
We first investigate the performance of the attention module using SP-Unified individually. We conducted a benchmark on two GPU nodes to measure the throughput of parallel Attention computations. We selected the same model parameter configuration as used in llama3-8B. Our code is publicly available at github repositoryhttps://github.com/feifeibear/long-context-attention.
As shown in Table 3, on an 8xA100-SXM4 NVLink node, the highest throughput for both 32K and 128K sequence lengths was achieved when the ulysses-degree is set to 8, that is the same as SP-Ulysses. Ulysses demonstrated a significant advantage over SP-Ring. The lb-ring indicates the load balance Ring-Attention. It consistently outperforms the basic ring attention implementation in terms of efficiency, with the advantage becoming more pronounced as the sequence length increases. This verified the argument in Sec. 3 of SP-Ring, that although communication overhead can be hidden through overlapping with computation, it results in a reduction in computational efficiency.
Table 3: Throughput (iters/sec) of SP-Unified on 8xA100-SXM4 NVLink fwd-only (Ring-Degree x Ulysses-Degree=8)
4 1 32K 32 128 8 135.569 136.375
4 1 32K 32 128 4 103.525 132.979
4 1 32K 32 128 2 91.365 132.979
4 1 32K 32 128 1 81.985 113.79
4 1 128K 32 128 8 2.782 2.785
4 1 128K 32 128 4 2.024 2.771
4 1 128K 32 128 2 1.73 2.89
4 1 128K 32 128 1 1.628 2.91
As shown in Table 4, on an 8xL20 PCIe node, the highest throughput for both 32K and 128K sequence lengths was achieved when the ulysses_degree is set to 4, when ring_degree=2.
Table 4: Throughput (iters/sec) of SP-Unified on 8xL20 PCIe fwd-only (Ring-Degree x Ulysses-Degree=8)
@@ 5.2 End-to-end SP Performance in Megatron-LM
We have integrated the SP-Unified method into Megatron-LM. As of now, Megatron-LM has an initial version of SP-Ring, but there is no implementation of Ulysses available. We did not utilize the open-source implementation of SP-Ulysses in Megatron-DeepSpeed, as we found that the performance of Megatron-DeepSpeed significantly lagged behind that of Megatron-LM. Additionally, we also discovered that the SP-Ulysses in Megatron-DeepSpeed was not compatible with Tensor Parallelism.
Our software is built on Megatron-LM commit ID 2196398, dated April 12, 2024. The SP-Ring implementation is based on Megatron-LM’s built-in Context Parallel, while SP-Ulysses is implemented using the code from our repository. It has already incorporated the load-balancing strategy discussed in this paper. We used the docker image nvcr.io/nvidia/nemo:24.03 for our experiments. We use ZeRO-1 for DP and SP by default. We always employ Sequence Parallelism for Tensor Parallelism, a.k.a TP-sp. We do not use gradient accumulation.
Our experimental hardware setup consists of two GPU nodes, each with 8xA800 GPUs connected via 400GB/s NVLink, and a node-to-node communication bandwidth of 1.6 Tbps RDMA.
@@
Firstly, we compare the performance of SP and DP under the same LLAMA2-7B workload on a single node of A800 GPUs. The global batch size is 8. We use the SP-Unified, and pick the best ulysses-degree and ring-degree settings. In a single node, ulysses-degree as 8 usually achieves the best performance. As shown in Figure 5, DP outperforms SP-Unified across various input sequence lengths, which confirms our conclusion in Tip 2.
@@ 5.4 Hybrid SP and TP
Table 5: LLAMA2-7B FLOPS per GPU of hybrid parallelism using TP and SP-Unified on 1 8xA800 NVLink node.
@@
Table 6: LLAMA3-8B FLOPS per GPU of hybrid parallelism using TP and SP-Unified on 2 8xA800 NVLink nodes.
@@
For sequence length are 64K and 80K, we can increase the global batch size (global-bs) for TP+SP hybrid setting when tp-degree=8. However, SP-only always has OOM issue with global-bs=2. Increasing the global-bs to 2, TP+SP has 2.7% improvement in throughput to SP-only at the sequence length 64K. But at sequence length 80K, TP+SP with global-bs as 2 is still worse than SP-only with global-bs as 1.
When the sequence length reaches 120K, the optimal performance is achieved with tp-degree=4 and ulysses-degree=2, reaching 266 TFLOPS, with an MFU of 0.85. At this sequence length, SP-only meets an OOM issue. This confirms our Tip 5 again, which implies that the memory efficiency of SP is inferior to that of TP-sp. It is noteworthy that two other TP+SP hybrid settings yield similar performance, 0.82 and 0.84 in MFU, which is quite close to the optimal.
@@
160K 1 4 2 2 284.5 0.91
160K 1 8 1 2 280.9 0.90
208K 1 8 1 2 269.8 0.86
160K 1 4 1 4 285.8 0.92
190K 1 4 1 4 285.8 0.92
We explored the upper bound of sequence length on 2 nodes, and the results are presented in Table 7. The largest sequence length was achieved with tp-degree=8 and ring-degree=2. At this point, SP-only could not run due to OOM. Compared to SP, TP-sp is more memory-efficient, hence for training the longest sequences, parallel degrees are limited by 8 attention head numbers should all be assigned to TP-sp.
As the sequence length increases from 64K to 208K, MFU is generally improving. This is due to the $O(L^{2})$ computational cost of attention quadratic increases, which results in a decreasing communication-to-computation ratio.
@@ 6 Future Work
Firstly, SP can introduce the dimension that can be partitioned from the sequence length, alleviating the constraints hindered by batch size limitations. MegaScale project [17] uses a large global batch size to increase DP degree which impacts convergence. However, we can increase the SP degree instead of DP degree to reduce the global batch size, thereby avoiding the convergence problem.

### v4

# USP: A Unified Sequence Parallelism Approach for Long Context Generative AI

@@ Abstract
Sequence parallelism (SP), which divides the sequence dimension of input tensors across multiple computational devices, is becoming key to unlocking the long-context capabilities of generative AI models. This paper investigates the state-of-the-art SP approaches, i.e. DeepSpeed-Ulysses and Ring-Attention, and proposes a unified SP approach, which is more robust to transformer model architectures and network hardware topology. This paper compares the communication and memory cost of SP and existing parallelism, including data/tensor/zero/pipeline parallelism, and discusses the best practices for designing hybrid 4D parallelism involving SP. We achieved 47% MFU on two 8xA800 nodes using SP for the LLAMA3-8B model training using sequence length 208K. Our code is publicly available at https://github.com/feifeibear/long-context-attention.
@@ 2 Sequence Parallelism Approaches
$L$ Sequence Length $d$ Hidden Dimension $hc$ Head Count
$hs$ Head Size $bs$ Batch Size $N$ Device Number
@@
The maturity of the standalone SP technology is marked by the publication of two milestone papers in late of 2023. The DeepSpeed-Ulysses [1], named as SP-Ulysses, and Ring-Attention [2], named as SP-Ring, solving the longstanding memory and communication issues inherent to SP from two distinct perspectives. For both methodologies, each computational device is allocated a distinct segment of the $Q$ (query), $K$ (key), $V$ (value), and $O$ (output) tensors, which are segregated along the dimensions of the sequence. There is no redundancy in their storage across devices, which is a primary distinction from the early SP design [6].
@@ 3 Unified Ulysses-Ring Sequence Parallelism
The inputs to the USP-Attention include the segment of the $Q,K,andV$ tensors after being partitioned along the sequence dimension and the output is a tensor $O$ shard. The size of a tensor segment is ($bs$, $L/N$, $hs$, $hd$). Note that when using MAQ, the shape of the $hc$ for the $K$ and $V$ tensor segments differs from that of the Q tensor. During forward propagation, $scatter\_idx$ is set to 1 and $gather\_idx$ is set to 2. The AllToAll4D merges the $L$ dimension and partitions the $hc$ dimension of the $Q$, $K$, and $V$ tensors, transforming them into ($hc/N$, $bs$, $L$, $d$). They also partition the $O$ tensor along the $L$ dimension and merge the $hc$ dimension. During the backward propagation, $scatter\_idx$ is set to 2, and $gather\_idx$ is set to 1.
Algorithm 1 Unified Sequence Parallelism Attention Implementation
1:function USP-Attn($ulysses\_pg$, $ring\_pg$, $Q$, $K$, $V$, $scatter\_idx$, $gather\_idx$)
2:     $Q\leftarrow\textbf{AllToAll4D}{(Q,scatter\_idx,gather\_idx,group=ulysses\_pg)}$
3:     $K\leftarrow\textbf{AllToAll4D}{(K,scatter\_idx,gather\_idx,group=ulysses\_pg)}$
4:     $V\leftarrow\textbf{AllToAll4D}{(V,scatter\_idx,gather\_idx,group=ulysses\_pg)}$
@@
6:     $O\leftarrow\textbf{AllToAll4D}{(O,gather\_idx,scatter\_idx,group=ulysses\_pg)}$
@@
Figure 3: Loading Balancing for SP-Ring.
The solution to the load-unbalancing issue is to reorder the input sequence tokens along the sequence dimension, as depicted on the right side of Figure 3. In the figure, the input sequence consists of 16 tokens. Under even partition, GPU0 processes tokens 0-3, while GPU3 handles tokens 12-15. After the reorder operation for load balance partitioning, GPU0 now processes tokens 0, 1, 14, 15, and GPU3 processes tokens 5,6,11,12. The workloads handled by each GPU are perfectly balanced, which is a superior solution to the striped attention [12].
The aforementioned load balancing method is quite straightforward and a similar implementation has already been applied in Megatron-LM. We present its application for Unified SP method. The Algorithm 2 extracts and reorders the input sequence from the global input sequence in hybrid Ulysses and Ring parallelism. For the positional encoding only involves element-wise operations, i.e. RoPE [13], the same reordering operation also needs to be applied to the positional encoding parameters of the model. Since the operation is applied to the model’s input token sequence, which is an integer vector of length $bs*L$, the additional overhead for load balancing is negligible.
@@
7:     $local\_seq\leftarrow reorder\_seq.chunk(ud)[u\_rank]$
@@
Firstly, Unified SP can remove the head number limitation of SP-Ulysses. For example, to run llama3-8B  [^3] https://github.com/meta-llama/llama3 with $hc$=8 in a 16-degree SP, one can set the Ulysses degree to 8 and the ring degree to 2. Secondly, Unified SP allows lower bandwidth and topology requirements for network infrastructure by offering a more robust communication pattern. By setting the Ulysses degree to a value between 1 and N, the Attention communication pattern will be a mix of P2P and All2All as shown in Figure 4. Such a communication pattern is particularly well-suited for heterogeneous communication networks, allowing All2All operations to operate in high-bandwidth interconnections while asynchronous P2P communications operate in lower-bandwidth sections. This is applicable to scenarios such as a node in which GPUs are connected via PCIe Switch or a cluster of GPU nodes between which the network is Ethernet connected.
@@
Figure 4: The Unified-SP is more robust to network hardware topology.
@@ 4 SP in 4D Parallelism
In terms of communication performance, SP is inferior to DP. Both SP and DP require the allreduce operation on gradients during backward propagation. The difference in their communication performance lies in the attention module, where SP introduces additional communications overhead for activations. When the Ulysses degree is set to be greater than 1, the communication overhead of SP will be larger than that of DP due to the all2all operations. When using the Ring method, although the additional P2P communication for attention is overlapped, it introduces extra performance issues. The ideal performance is only to reach the performance of DP without communication for attention. In terms of memory performance, both SP and DP are equivalent, as they all can reduce the activation footprint to 1/$N$.
@@
ZeRO [14] is a distributed parameter management method that reduces the storage space requirements of each computing device by sharding the Optimizer States (ZeRO-1), Gradients (ZeRO-2), and Parameters (ZeRO-3) across multiple devices. The memory cost for Optimizer States, Gradients, and Parameters is reduced to 1/$N$ of the original. ZeRO can also operate within an SP process group because partitioning along the batch dimension ($bs$) or the sequence dimension ($L$) is equivalent to ZeRO’s approach. For ZeRO, SP and DP process groups can be viewed as a combined process group of size $N_{s}p\times N_{d}p$.
@@
PP partitions transformer blocks across layers, so it is complementary with the SP, which partitions tensors inside a transformer block. Therefore, SP and PP are fully compatible. Since SP can form a unified parallel group with ZeRO for collective communication, we believe TP should still be placed at the lowest dimension of the 4D parallel group partitioning.
@@ 5.1 Performance of Unified SP
We evaluated the attention module’s performance using SP-Unified on an L20 PCIe GPU cluster, benchmarking the throughput in TFLOPS. We replicated the parameter settings of the llama3-8B model. Table 3 illustrates that on an 8xL20 setup, optimal throughput for 32K and 128K sequences was achieved with a ulysses-degree of 4 and ring-degree set to 2. This finding supports Tip 1: SP-Unified is well-suited for heterogeneous networks. The lb-ring, an enhanced version of the standard Ring-Attention with load-balancing, outperformed the original. It consistently outperforms the basic ring attention implementation, with the advantage becoming more pronounced as the sequence length increases.
Table 3: Throughput (iters/sec) of SP-Unified on 8xL20 PCIe fwd-only (Ring-Degree $\times$ Ulysses-Degree=8)
@@
As shown in Table 4, we repeated the attention benchmarking on an 8xA100-SXM4 NVLink node, the highest throughput for both 32K and 128K sequence lengths was achieved when the ulysses-degree is set to 8, that is the same as SP-Ulysses. SP-Ulysses demonstrated a significant advantage over SP-Ring. This verified the argument in Sec. 3 of SP-Ring, that although communication overhead can be hidden through overlapping with computation, it results in a reduction in computational efficiency.
Table 4: Throughput (iters/sec) of SP-Unified on 8xA100-SXM4 NVLink fwd-only (Ring-Degree $\times$ Ulysses-Degree=8)
4 1 32K 32 128 8 135.569 136.375
4 1 32K 32 128 4 103.525 132.979
4 1 32K 32 128 2 91.365 132.979
4 1 32K 32 128 1 81.985 113.79
4 1 128K 32 128 8 2.782 2.785
4 1 128K 32 128 4 2.024 2.771
4 1 128K 32 128 2 1.73 2.89
4 1 128K 32 128 1 1.628 2.91
@@ 5.2 End-to-end SP Performance in Megatron-LM
We have incorporated the SP-Unified method into Megatron-LM. Currently, Megatron-LM includes a preliminary version of SP-Ring, but lacks an implementation of SP-Ulysses. Our software is based on Megatron-LM commit 2196398, dated April 12, 2024. The SP-Ring is implemented using Megatron-LM’s native Context Parallel, while SP-Ulysses is developed from our repository code. We conducted experiments using the docker image nvcr.io/nvidia/nemo:24.03. By default, we utilize ZeRO-1 for both Data Parallelism (DP) and Sequence Parallelism (SP), and consistently apply Sequence Parallelism Optimization for Tensor Parallelism (TP-sp). We do not employ gradient accumulation or activation recomputation. Our experimental setup includes two GPU nodes, each equipped with 8xA800 GPUs connected via 400GB/s NVLink, and node-to-node communication via 1.6 Tbps RDMA. We have adjusted the MFU computation in Megatron-LM to account only for effective computations under causal masking. As a result, the MFU is notably lower when training with long sequences compared to the MFU figures printed by Megatron-LM.
@@
Firstly, we compare the performance of SP and DP under the same LLAMA2-7B workload on a single node of A800 GPUs. The global batch size is 8. We use the SP-Unified, and pick the best ulysses-degree and ring-degree settings. In a single node, ulysses-degree = 8 usually achieves the best performance. As shown in Figure 5, DP outperforms SP-Unified across various input sequence lengths, which confirms our conclusion in Tip 2.
@@ 5.4 Hybrid SP and TP
Table 5: LLAMA2-7B FLOPS per GPU of hybrid parallelism using TP and SP-Unified on a single 8xA800 NVLink node.
@@
Table 6: LLAMA3-8B FLOPS per GPU of hybrid parallelism using TP and SP-Unified on two RDMA-connected 8xA800 NVLink nodes.
@@
For sequence lengths are 64K and 80K, we can increase the global batch size (global-bs) for the TP+SP hybrid setting when tp-degree=8. However, SP-only always has an OOM issue with global-bs=2. Increasing the global-bs to 2, TP+SP has a 2.7% improvement in throughput to SP-only at the sequence length 64K. But at sequence length 80K, TP+SP with global-bs as 2 is still worse than SP-only with global-bs as 1.
When the sequence length reaches 120K, the optimal performance is achieved with tp-degree=4 and ulysses-degree=2, reaching 152.51 TFLOPS, with an MFU of 0.49. At this sequence length, SP-only meets an OOM issue. This confirms our Tip 5 again, which implies that the memory efficiency of SP is inferior to that of TP-sp. It is noteworthy that two other TP+SP hybrid settings yield similar performance, 0.47 and 0.48 in MFU, which is quite close to the optimal.
@@
160K 1 4 2 2 158.64 0.51
160K 1 8 1 2 156.63 0.50
208K 1 8 1 2 147.26 0.47
160K 1 4 1 4 159.37 0.51
190K 1 4 1 4 157.08 0.50
We explored the upper bound of sequence length on 2 nodes, and the results are presented in Table 7. The largest sequence length was achieved with tp-degree=8 and ring-degree=2. At this point, SP-only could not run due to OOM. Compared to SP, TP-sp is more memory-efficient, hence for training the longest sequences, parallel degrees are limited by 8 attention head numbers should all be assigned to TP-sp. It is noteworthy that we can achieve longer sequence lengths through activation optimization, such as activation recomputation [16], and offloading [5, 15], but these methods will harm MFU.
@@ 6 Future Work
Firstly, SP can introduce the dimension that can be partitioned from the sequence length, alleviating the constraints hindered by batch size limitations. MegaScale project [18] uses a large global batch size to increase DP degree which impacts convergence. However, we can increase the SP degree instead of the DP degree to reduce the global batch size, thereby avoiding the convergence problem.